40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Fri Jul 16 23:28:24 2021, Koji, Summary, Cryo vacuum chamber, Sus Test Work 07/16/2021 7x
    Reply  Sat Jul 17 00:55:41 2021, Koji, Summary, Cryo vacuum chamber, Temp Log 210716_2255 temp_log_cool_down_20210716_2255.pdf
       Reply  Mon Jul 19 15:57:17 2021, Stephen, Summary, Cryo vacuum chamber, Temp Log 210716_2255 
          Reply  Mon Jul 19 17:21:19 2021, Koji, Summary, Cryo vacuum chamber, Temp Log 210716_2255 temp_log_cool_down_20210716_2255.pdf
             Reply  Tue Jul 20 11:33:52 2021, Koji, Summary, Cryo vacuum chamber, A cooling model (Temp Log 210716_2255) cooling_model.pdfIMG_9390.JPGIMG_9391.JPG
                Reply  Tue Jul 20 17:28:30 2021, Koji, Summary, Cryo vacuum chamber, A cooling model (Temp Log 210716_2255) cooling_model.pdf
                   Reply  Wed Jul 21 21:05:59 2021, Koji, Summary, Cryo vacuum chamber, Test mass cooling (2021/07/16 ~ 2021/07/21) oplev_trend.png20210721201333_IMG_0765.jpeg20210716234113_IMG_0742.jpegScreenshot_from_2021-07-21_20-19-09.png
                      Reply  Thu Jul 22 22:03:45 2021, Koji, Summary, Cryo vacuum chamber, Test mass heating in progress (2021/07/21 ~ 2021/07/23) IMG-9395.jpgtemp_log_warmup_20210721_2052.pdf
                      Reply  Sun Jul 25 21:45:46 2021, Koji, Summary, Cryo vacuum chamber, About the radiation heat transfer model 
Message ID: 2614     Entry time: Wed Jul 21 21:05:59 2021     In reply to: 2611     Reply to this: 2615   2617
Author: Koji 
Type: Summary 
Category: Cryo vacuum chamber 
Subject: Test mass cooling (2021/07/16 ~ 2021/07/21) 

[Stephen and Koji for discussion / Koji for the execution]

1. Temperature Trend

See [QIL ELOG 2611] for the updated temp log and the cooling model.

Considerations for the next cycle:
-> How can we accelerate the cooling? It seems that the table cooling is conduction limited. Improve the cold head connection.
-> We want to move the RDTs
-> How can we improve radiative cooling?

2. Oplev Trend (Attachment 1)

Sum: The beam has been always on the QPD (good). See also Attachment 2

X&Y: In the first few hours the beam drifted in -X and then +X while Y had slow continuous drift in +Y. ~11hours later sudden drift in -Y and totally saturated. Also -X saturation observed @~16hrs. Again +Y drift was seen @~25hrs. The totally saturated in -X and +Y.
They may be related to the drift of various components with various cooling time scale.

Visual check: ~2mm shift in X&Y is visually observed. Attachment 2

-> How can we quantify the drift? What information do we want to extract?

3. OSEM and the magnet

The magnet is intact. And the suspension seemed still free after cooling (Attachment 3)
Significant misalignment was not visible. No visible damage by cooling was found. The coil is alive and the PD/LED are also intact. Fluke showed that they are still diodes, but their function was not checked.

The coil resistance changed from 16Ohm -> 4.2Ohm. For the 16Ohm, 2 Ohm was from the wire. Let's assume we still have 2Ohm overhead -> The coil R changed from 14->2.2. This corresponds to the coil temperature of the order of ~100K. This is not so crazy.

Some actuation current was applied to the magnet. For this test, the oplev was realigned.
First, some ~300mA current pulses were applied to the coil. The ringdown of the yaw mode was visible. Then the DC current of 100mA was applied. This didn't make visible change on the spot position but the data showed that there was a DC shift.

-> We prefer to have a softer suspension for the next test.

4. CTC100 logging

During the cooling we kept having inaccurate data logged compared with the displayed data on the screen of CTC100.
As soon as the cooling logging was stopped, telneting to CTC100 was available. So, I telnetted to the device and sent the data transfer command ("getOutput"). Surprisingly, the returned values agreed with the displayed values.
So my hypothesis is that somehow the data strings are buffered somewhere and gradually the returned values get delayed. From the behavior of the device, I imagined that the fresh telnet connection gives us the latest data and there is no buffering issue.

So I tweaked the data logging code to establish the telnet connection every time the values are asked. The connection is closed after the every data acquisition. I like this as we can also make the test connection between each data acquisition points, although I have not tried it yet. The code is in the same folder named ctc100_controller_v2.py

5. Heating

Now I thought that I did all I wanted to do this evening, so the heater was turned on at ~20:50, Jul 21. The heating power saturated at 22W, which is the set limit.

Attachment 1: oplev_trend.png  51 kB  Uploaded Wed Jul 21 22:40:06 2021  | Hide | Hide all
oplev_trend.png
Attachment 2: 20210721201333_IMG_0765.jpeg  961 kB  Uploaded Wed Jul 21 22:40:19 2021  | Hide | Hide all
20210721201333_IMG_0765.jpeg
Attachment 3: 20210716234113_IMG_0742.jpeg  1.782 MB  Uploaded Wed Jul 21 22:42:44 2021  | Hide | Hide all
20210716234113_IMG_0742.jpeg
Attachment 4: Screenshot_from_2021-07-21_20-19-09.png  116 kB  Uploaded Wed Jul 21 22:47:48 2021  | Hide | Hide all
Screenshot_from_2021-07-21_20-19-09.png
ELOG V3.1.3-